# Crop Pest Classification Using Deep Learning Techniques: A Review

### v1

@@ 4.6 Other Crops
Ramcharan et al.[4] Cassava CNN (ResNet-50, transfer) 5,660 field images (5 classes) $\sim$93acc (field) Field conditions cause some errors (lighting, background).
@@
Nagalakshmi et al. [16] Peanut Hybrid CNN+ViT (EViTA + MFO) Peanut pest images ($\sim$2k, 4 classes) $\sim$92acc, 0.95 precision Improved over CNN; custom dataset, specific to region.
Venkatasaichandrakanth [15] Peanut Hybrid (dual-transformer + MFO) Peanut pest images ($\sim$2k, similar to above) $\sim$92acc, high specificity Similar approach to [16] confirms transformer benefit.
@@ 5.3 Other Insects
{adjustwidth}
-2.25 in0in Ref Pest Type Examples (Species) Datasets / Crops Remarks [17], [7], [13] Sap-Sucking Insects Aphids, Whiteflies, Leafhoppers, Scales IP102 (many aphid/whitefly classes); Tea pest images (custom); CactiViT dataset (cactus) Tiny size; require high-res and attention. Class imbalance common (few samples of certain species). [26],[2],[13] Caterpillars & Borers Fall armyworm, Stem borers, Leaf miner larvae PlantVillage (includes some pest damage); Custom field images (maize, tomato) Larger larvae easier to classify if visible. Hidden borers still a challenge (may rely on damage signs). [27],[28],[29] Beetles & Weevils Grain weevils, Colorado potato beetle, Palm weevil Xie et al. 2018 (40 pest dataset); NBAIR (India) pest images; Custom trap images Many lab datasets exist (isolated insect images). Field detection requires dealing with camouflage and varied backgrounds. [30], [31] Other (Flies, etc.) Fruit flies, Locusts, Thrips, Mites, etc. Custom trap image sets; Drone imagery (locusts); Mixed insect image datasets Highly variable group. Large-scale models struggle across all (e.g., IP102). Few-shot learning emerging to handle rare pests.
@@ 6.1 CNN Based Approaches
{adjustwidth}
-2.25 in0in Ref Year Description [32] 2020 Evaluated 7 CNN architectures (VGG16, ResNet50, InceptionV3, etc.) for multi-pest classification; top models achieved $\sim$97accuracy. [3] 2025 Used GoogleNet (CNN) as a baseline for pest ID in Jordan; highlighted CNNs’ strong performance on local pest datasets. [33] 2023 Employed EfficientNetV2 CNN to handle high variability in field images, improving pest classification accuracy. [23] 2024 Proposed GA-GhostNet (a lightweight CNN with attention) achieving competitive accuracy with dramatically fewer parameters. [34] 2023 Applied CNN models to X-ray images of fruit fly pupae; achieved $>$97% accuracy in classifying parasitized vs. healthy pupae.
@@ 6.2 Vision Transformers
{adjustwidth}
-2.25 in0in Reference Year Description [5] 2021 Introduced Vision Transformer (ViT) for image recognition, using self-attention on image patches instead of convolutions, a foundational work enabling transformers in vision. [22] 2022 Applied a ViT to pest images with extensive augmentation; achieved $\sim$96accuracy, outperforming CNN baselines on the same pest/disease dataset. [35] 2024 Improved ViT’s focus via block-wise image partitioning, helping the model attend to key pest regions (e.g., insect body), which boosted classification accuracy. [25] 2023 Embedded a Swin Transformer into YOLO (YOLO-GBS) for rice pest detection; captured long-range dependencies in cluttered field images, improving detection performance. [12] 2025 Developed a fully transformer-based pest detector with RPSA attention; excelled at modeling global context and detecting small pests in images.
@@ 6.3 Hybrid CNN-Transformer Architectures
{adjustwidth}
-2.25in0in Reference Year Description [16] 2024 Developed EViTA, a hybrid model with CNN feature extractor + transformer encoder, outperforming standalone CNN or ViT on peanut pest images. [15] 2023 Extended EViTA architecture; integrated Moth Flame Optimization (MFO) for feature selection and parameter tuning, further enhancing pest classification precision. [36] 2023 Hybrid CNN-transformer for tea pests: CNN layers for local features, followed by transformer blocks for global context—improved performance in cluttered field images. [37] 2024 Embedded a hybrid pooled multi-head attention module into ViT, fusing CNN pooling with self-attention; achieved higher accuracy on fine-grained pest recognition. [38] 2025 Introduced ConvViT for insect detection, a CNN+ViT hybrid optimized for mobile deployment, combining high accuracy with low model size (edge-friendly). [11] 2024 Proposed Pest-ConFormer, a multi-scale hybrid CNN-transformer architecture for large-scale pest recognition, achieving state-of-the-art results on a 14-class pest dataset.
@@ 6.4 Object Detection Models
{adjustwidth}
-2.25in0in Reference Year Description [10] 2023 Employed YOLOv5 for pest localization in field images; demonstrated one-stage detectors can accurately detect multiple pests per image. [13] 2024 Applied YOLOv7 to pest detection with improved accuracy; utilized hard example mining to better learn from challenging, camouflaged pest instances. [6] 2025 Deployed YOLOv8 in a pest monitoring system; achieved $R^{2}=0.99$ for automated pest counts, validating detection models for precise population estimation. [9] 2023 Systematic review of insect detection; noted YOLO models consistently outperform Faster R-CNN on pest datasets in both accuracy and speed. [40] 2023 Developed a lightweight YOLOv5-based detector for rice pests; added a large-scale feature layer to improve detection of small pests, boosting recall significantly. [39] 2023 Created “Insect-YOLO” with custom enhancements for multi-crop pest detection (streamlined parameters, specialized augmentations); achieved real-time detection with high accuracy.
@@ 6.5 Emerging and Non-Traditional AI Techniques
{adjustwidth}
-2.25in0in Reference Year Description [10] 2023 Explored a quantum convolutional neural network for pest classification; achieved near-perfect accuracy on a small dataset, showing early promise for quantum-enhanced learning. [24] 2022 Used self-supervised BYOL to pre-train on unlabeled pest images; fine-tuned model reached $\sim$94accuracy on 12-class pest identification, nearly matching fully supervised results. [16] 2024 Incorporated Moth Flame Optimization (MFO) in their hybrid model’s training; this metaheuristic optimized feature selection and hyperparameters, boosting classification performance. [41] 2022 Applied a diffusion model to generate synthetic pest images for data augmentation; improved long-tail pest class accuracy significantly by augmenting scarce classes. [42] 2022 Developed a GAN-based augmentation tool for pest detection (case study: whiteflies on traps); augmented data led to higher detection rates for rare pest instances. [43] 2022 Integrated pest CNN models with an Augmented Reality app for farmers; enabled real-time pest identification through smartphone camera with overlaid guidance (enhancing practical usability).
@@
The technical landscape of AI in pest classification is rapidly evolving from foundational CNN classifiers to transformer-centric and hybrid architectures, and even experimental paradigms like quantum learning and self-supervised approaches. The literature consistently shows that using advanced architectures such as ViTs, attention modules, or hybrid combinations yields significant performance improvements. How- ever, such gains are often balanced against increased model complexity and resource demands. Authors like Saranya et al.[37] have explicitly noted the trade-offs between accuracy and deployability. As a result, there is growing interest in designing lightweight yet effective models. For example, Xiao et al.[23] and Wang et al.[6] focused on efficiency through model compression or architectural streamlining to enable on- device deployment an important consideration for field applications where compute resources may be limited. Ultimately, the diversity of approaches provides researchers and practitioners with a robust toolkit. The choice of model architecture and training strategy will depend heavily on specific use-case requirements, such as the need for speed, interpretability, dataset size, or deployment environment. This rich ecosystem of AI techniques underscores an exciting frontier in pest management, where technology is increasingly capable of providing timely, accurate, and scalable solutions.Figure 3 demonstrates the usage frequency of AI techniques in pest classification.
@@ 7.5 Other Datasets
{adjustwidth}
-2.25in0in Ref Dataset (Year) Classes Images Public Typical Image Resolution Notes [8] IP102 (2019) 102 $\sim$75,000 Yes 500$\times$500 px (varied; many field shots) Benchmark for fine-grained pest recognition. [44] Xie1 (2015) 24 1,440 Semi $\sim$600$\times$600 px (est., lab images) Early dataset of common pest; small per class (60). Often used for initial model tests. [45] Xie2/D0 (2018) $\sim$40–45 $\sim$4,500 Semi $\sim$800$\times$800 px (est.) Includes more species (possibly storage pests). [46] PlantVillage-39 (2016) 39 (mixed) $\sim$54,000 (subset) Yes 256$\times$256 px (processed) Lab-like images, high quality. [3] Tomato Pest (2025) 2 (infested vs healthy) 859 (raw) / 5,894 (aug) No 1080$\times$1080 px (smartphone) Augmented for training. Focused binary classification. [7] CactiViT (2023) 3 (infestation levels) $\sim$1,200 Yes 720$\times$1280 px (smartphone) Open-source on GitHub; domain-specific. [48] Kaggle Paddy Pest (2022) 2 (pest vs none) 648 Yes 640$\times$480 px (varied) Useful for simple detection; highly unbalanced. [47] Rice Pest 12 (2016) 12 5,136 No $\sim$400$\times$400 to 800$\times$800 px (field) Rice field pest image. Moderately sized, used in early CNN work. [13] Corn-Rice Pest (2024) $\sim$20 (est.) few thousand (est.) No 1024$\times$1024 px (drone/field photos) Du et al.’s YOLOv7-PSAFP data. Focused on small pests in field. Possibly will be released in future. [12] Forestry Pest (2025) 15–30 (est.) $\sim$10,000 (est.) No 1280$\times$720 px (field) Liu et al.’s transformer dataset. Uniform background (tree bark/leaves) helped high mAP.
@@ 8.1 Dataset Imbalance
{adjustwidth}
-2.25in0in Reference Year Description [35] 2024 Noted significant accuracy drop on under-represented pest classes; highlighted how long-tailed training data skewed ViT performance. [15] 2024 Applied focal loss to reduce bias toward frequent classes in pest classification, giving more weight to minority class examples. [3] 2025 Emphasized the critical shortage of localized pest image datasets (e.g., for Jordan crops), identifying data scarcity as a key limitation. [20] 2023 Similarly pointed out the lack of comprehensive, annotated pest image repositories; called for community efforts in data collection and sharing. [24] 2022 Demonstrated self-supervised pre-training (BYOL) on unlabeled pest images to combat data scarcity; achieved near supervised-level accuracy with far fewer labeled examples. [8] 2019 Released IP102, a 75,000+ image dataset of 102 pest species with a naturally imbalanced distribution, a benchmark that spurred research into long-tail learning for pests. [49] 2021 Created AgriPest, a large-scale pest dataset ($\sim$49.7k images, 14 species) captured in real field conditions; aims to provide a realistic benchmark for training and evaluation.
@@ 8.2 Complex Field Environments
{adjustwidth}
-2.25in0in Reference Year Description [12] 2025 Introduced a Region-based Pyramid Self-Attention in their detector to handle cluttered backgrounds; improved pest detection in visually busy field images by focusing on relevant regions. [13] 2024 Employed hard example mining during training to teach the model using cluttered, difficult images; reduced false positives from leaves and background noise. [17] 2025 Observed model accuracy degrading with rain, fog, or harsh shadows; highlighted need for augmentation simulating weather and lighting variability for robust performance. [41] 2022 Developed a deformable residual network with a global context module for pest detection in complex scenes; achieved 77.8% mAP on 21 pest classes in field conditions by explicitly modeling context. [39] 2023 Proposed environment-invariant training techniques (e.g., multi-domain data and adaptive normalization) enabling their Insect-YOLO to maintain accuracy across different backgrounds and lighting conditions.
@@ 8.3 Detection of Small Pests
{adjustwidth}
-2.25in0in Reference Year Description [25] 2023 Added an extra fine-scale detection head in their YOLO-based model to capture very small pests; improved detection of tiny insects that standard heads missed. [6] 2025 Incorporated CBAM attention into the detection pipeline, heightening the model’s sensitivity to small, localized pest features (e.g., aphids on a leaf). [9] 2023 Identified small object detection as a major challenge across studies; noted that models perform significantly worse on insects that occupy few pixels, urging focus on multi-scale methods. [50] 2024 Developed a method for detecting whiteflies on yellow sticky traps (small targets); used high-resolution input and an improved YOLOv5, achieving high precision on pests $\sim$3mm in size. [51] 2023 Applied image super-resolution on suspected pest regions before classification; demonstrated that upsampling tiny pest images improved identification accuracy in a lab setting.
@@ 8.4 Computational Cost
{adjustwidth}
-2.25in0in Reference Year Description [3] 2025 Emphasized mobile deployment requirements; aimed to develop models that run on smartphones for on-site pest identification. [7] 2023 Demonstrated a successful on-device pest classification app; used model optimization (quantization, etc.) to achieve real-time inference on mobile hardware. [23] 2024 Utilized GA-GhostNet (an efficient CNN) to balance accuracy and speed; significantly reduced model parameters while maintaining high precision. [37] 2024 Designed their hybrid model with edge constraints in mind (e.g., limiting extra parameters from HP-MHA module) to ensure feasibility of deployment. [52] 2022 Created a lightweight YOLOv3-based detector ($\sim$5M params) that runs $\sim$20 FPS on edge device; achieved $>$80% mAP, illustrating a strong accuracy-speed trade-off for field use. [53] 2024 Improved YOLOv8 architecture for efficiency (small dataset of 7 classes); obtained 82.3% mAP with only 2M parameters, suitable for low-power deployment. [54] 2015 Proposed knowledge distillation, a method yet to be fully explored in pest classification, where a smaller model learns from a larger model’s outputs, potentially yielding a compact, accurate pest classifier.

### v3

@@ 4.6 Other Crops
Ramcharan et al.[4] Cassava CNN (ResNet-50, transfer) 5,660 field images (5 classes) $\sim$93% acc (field) Field conditions cause some errors (lighting, background).
@@
Nagalakshmi et al. [16] Peanut Hybrid CNN+ViT (EViTA + MFO) Peanut pest images ($\sim$2k, 4 classes) $\sim$92% acc, 0.95 precision Improved over CNN; custom dataset, specific to region.
Venkatasaichandrakanth [15] Peanut Hybrid (dual-transformer + MFO) Peanut pest images ($\sim$2k, similar to above) $\sim$92% acc, high specificity Similar approach to [16] confirms transformer benefit.
@@ 5.3 Other Insects
Ref Pest Type Examples (Species) Datasets / Crops Remarks
[17], [7], [13] Sap-Sucking Insects Aphids, Whiteflies, Leafhoppers, Scales IP102 (many aphid/whitefly classes); Tea pest images (custom); CactiViT dataset (cactus) Tiny size; require high-res and attention. Class imbalance common (few samples of certain species).
[26],[2],[13] Caterpillars & Borers Fall armyworm, Stem borers, Leaf miner larvae PlantVillage (includes some pest damage); Custom field images (maize, tomato) Larger larvae easier to classify if visible. Hidden borers still a challenge (may rely on damage signs).
[27],[28],[29] Beetles & Weevils Grain weevils, Colorado potato beetle, Palm weevil Xie et al. 2018 (40 pest dataset); NBAIR (India) pest images; Custom trap images Many lab datasets exist (isolated insect images). Field detection requires dealing with camouflage and varied backgrounds.
[30], [31] Other (Flies, etc.) Fruit flies, Locusts, Thrips, Mites, etc. Custom trap image sets; Drone imagery (locusts); Mixed insect image datasets Highly variable group. Large-scale models struggle across all (e.g., IP102). Few-shot learning emerging to handle rare pests.
@@ 6.1 CNN Based Approaches
Ref Year Description
[32] 2020 Evaluated 7 CNN architectures (VGG16, ResNet50, InceptionV3, etc.) for multi-pest classification; top models achieved $\sim$97% accuracy.
[3] 2025 Used GoogleNet (CNN) as a baseline for pest ID in Jordan; highlighted CNNs’ strong performance on local pest datasets.
[33] 2023 Employed EfficientNetV2 CNN to handle high variability in field images, improving pest classification accuracy.
[23] 2024 Proposed GA-GhostNet (a lightweight CNN with attention) achieving competitive accuracy with dramatically fewer parameters.
[34] 2023 Applied CNN models to X-ray images of fruit fly pupae; achieved >>97% accuracy in classifying parasitized vs. healthy pupae.
@@ 6.2 Vision Transformers
Reference Year Description
[5] 2021 Introduced Vision Transformer (ViT) for image recognition, using self-attention on image patches instead of convolutions, a foundational work enabling transformers in vision.
[22] 2022 Applied a ViT to pest images with extensive augmentation; achieved $\sim$96% accuracy, outperforming CNN baselines on the same pest/disease dataset.
[35] 2024 Improved ViT’s focus via block-wise image partitioning, helping the model attend to key pest regions (e.g., insect body), which boosted classification accuracy.
[25] 2023 Embedded a Swin Transformer into YOLO (YOLO-GBS) for rice pest detection; captured long-range dependencies in cluttered field images, improving detection performance.
[12] 2025 Developed a fully transformer-based pest detector with RPSA attention; excelled at modeling global context and detecting small pests in images.
@@ 6.3 Hybrid CNN-Transformer Architectures
Reference Year Description
[16] 2024 Developed EViTA, a hybrid model with CNN feature extractor + transformer encoder, outperforming standalone CNN or ViT on peanut pest images.
[15] 2023 Extended EViTA architecture; integrated Moth Flame Optimization (MFO) for feature selection and parameter tuning, further enhancing pest classification precision.
[36] 2023 Hybrid CNN-transformer for tea pests: CNN layers for local features, followed by transformer blocks for global context—improved performance in cluttered field images.
[37] 2024 Embedded a hybrid pooled multi-head attention module into ViT, fusing CNN pooling with self-attention; achieved higher accuracy on fine-grained pest recognition.
[38] 2025 Introduced ConvViT for insect detection, a CNN+ViT hybrid optimized for mobile deployment, combining high accuracy with low model size (edge-friendly).
[11] 2024 Proposed Pest-ConFormer, a multi-scale hybrid CNN-transformer architecture for large-scale pest recognition, achieving state-of-the-art results on a 14-class pest dataset.
@@ 6.4 Object Detection Models
Reference Year Description
[10] 2023 Employed YOLOv5 for pest localization in field images; demonstrated one-stage detectors can accurately detect multiple pests per image.
[13] 2024 Applied YOLOv7 to pest detection with improved accuracy; utilized hard example mining to better learn from challenging, camouflaged pest instances.
[6] 2025 Deployed YOLOv8 in a pest monitoring system; achieved $R^{2}=0.99$ for automated pest counts, validating detection models for precise population estimation.
[9] 2023 Systematic review of insect detection; noted YOLO models consistently outperform Faster R-CNN on pest datasets in both accuracy and speed.
[40] 2023 Developed a lightweight YOLOv5-based detector for rice pests; added a large-scale feature layer to improve detection of small pests, boosting recall significantly.
[39] 2023 Created “Insect-YOLO” with custom enhancements for multi-crop pest detection (streamlined parameters, specialized augmentations); achieved real-time detection with high accuracy.
@@ 6.5 Emerging and Non-Traditional AI Techniques
Reference Year Description
[10] 2023 Explored a quantum convolutional neural network for pest classification; achieved near-perfect accuracy on a small dataset, showing early promise for quantum-enhanced learning.
[24] 2022 Used self-supervised BYOL to pre-train on unlabeled pest images; fine-tuned model reached $\sim$94% accuracy on 12-class pest identification, nearly matching fully supervised results.
[16] 2024 Incorporated Moth Flame Optimization (MFO) in their hybrid model’s training; this metaheuristic optimized feature selection and hyperparameters, boosting classification performance.
[41] 2022 Applied a diffusion model to generate synthetic pest images for data augmentation; improved long-tail pest class accuracy significantly by augmenting scarce classes.
[42] 2022 Developed a GAN-based augmentation tool for pest detection (case study: whiteflies on traps); augmented data led to higher detection rates for rare pest instances.
[43] 2022 Integrated pest CNN models with an Augmented Reality app for farmers; enabled real-time pest identification through smartphone camera with overlaid guidance (enhancing practical usability).
The technical landscape of AI in pest classification is rapidly evolving from foundational CNN classifiers to transformer-centric and hybrid architectures, and even experimental paradigms like quantum learning and self-supervised approaches. The literature consistently shows that using advanced architectures such as ViTs, attention modules, or hybrid combinations yields significant performance improvements. However, such gains are often balanced against increased model complexity and resource demands. Authors like Saranya et al.[37] have explicitly noted the trade-offs between accuracy and deployability. As a result, there is growing interest in designing lightweight yet effective models. For example, Xiao et al.[23] and Wang et al.[6] focused on efficiency through model compression or architectural streamlining to enable on-device deployment, an important consideration for field applications where compute resources may be limited. Ultimately, the diversity of approaches provides researchers and practitioners with a robust toolkit. The choice of model architecture and training strategy will depend heavily on the specific use-case requirements, such as the need for speed, interpretability, dataset size, or deployment environment. This rich ecosystem of AI techniques underscores an exciting frontier in pest management, where technology is increasingly capable of providing timely, accurate, and scalable solutions. Figure 3 demonstrates the usage frequency of AI techniques in pest classification.
@@ 7.5 Other Datasets
Ref Dataset (Year) Classes Images Public Typical Image Resolution Notes
[8] IP102 (2019) 102 $\sim$75,000 Yes $\times500$ px (varied; many field shots) Benchmark for fine-grained pest recognition.
[44] Xie1 (2015) 24 1,440 Semi $\sim$$\times600$ px (est., lab images) Early dataset of common pest; small per class (60). Often used for initial model tests.
[45] Xie2/D0 (2018) $\sim$40–45 $\sim$4,500 Semi $\sim$$\times800$ px (est.) Includes more species (possibly storage pests).
[46] PlantVillage-39 (2016) 39 (mixed) $\sim$54,000 (subset) Yes $\times256$ px (processed) Lab-like images, high quality.
[3] Tomato Pest (2025) 2 (infested vs healthy) 859 (raw) / 5,894 (aug) No $\times1080$ px (smartphone) Augmented for training. Focused binary classification.
[…]
[47] Rice Pest 12 (2016) 12 5,136 No $\sim$$\times400$ to $\times800$ px (field) Rice field pest image. Moderately sized, used in early CNN work.
[13] Corn-Rice Pest (2024) $\sim$20 (est.) few thousand (est.) No $\times1024$ px (drone/field photos) Du et al.’s YOLOv7-PSAFP data. Focused on small pests in field. Possibly will be released in future.
[…]
@@ 8.1 Dataset Imbalance
Reference Year Description
[35] 2024 Noted significant accuracy drop on under-represented pest classes; highlighted how long-tailed training data skewed ViT performance.
[15] 2024 Applied focal loss to reduce bias toward frequent classes in pest classification, giving more weight to minority class examples.
[3] 2025 Emphasized the critical shortage of localized pest image datasets (e.g., for Jordan crops), identifying data scarcity as a key limitation.
[20] 2023 Similarly pointed out the lack of comprehensive, annotated pest image repositories; called for community efforts in data collection and sharing.
[24] 2022 Demonstrated self-supervised pre-training (BYOL) on unlabeled pest images to combat data scarcity; achieved near supervised-level accuracy with far fewer labeled examples.
[8] 2019 Released IP102, a 75,000+ image dataset of 102 pest species with a naturally imbalanced distribution, a benchmark that spurred research into long-tail learning for pests.
[49] 2021 Created AgriPest, a large-scale pest dataset ($\sim$49.7k images, 14 species) captured in real field conditions; aims to provide a realistic benchmark for training and evaluation.
@@ 8.2 Complex Field Environments
Reference Year Description
[12] 2025 Introduced a Region-based Pyramid Self-Attention in their detector to handle cluttered backgrounds; improved pest detection in visually busy field images by focusing on relevant regions.
[13] 2024 Employed hard example mining during training to teach the model using cluttered, difficult images; reduced false positives from leaves and background noise.
[17] 2025 Observed model accuracy degrading with rain, fog, or harsh shadows; highlighted need for augmentation simulating weather and lighting variability for robust performance.
[41] 2022 Developed a deformable residual network with a global context module for pest detection in complex scenes; achieved 77.8% mAP on 21 pest classes in field conditions by explicitly modeling context.
[39] 2023 Proposed environment-invariant training techniques (e.g., multi-domain data and adaptive normalization) enabling their Insect-YOLO to maintain accuracy across different backgrounds and lighting conditions.
@@ 8.3 Detection of Small Pests
Reference Year Description
[25] 2023 Added an extra fine-scale detection head in their YOLO-based model to capture very small pests; improved detection of tiny insects that standard heads missed.
[6] 2025 Incorporated CBAM attention into the detection pipeline, heightening the model’s sensitivity to small, localized pest features (e.g., aphids on a leaf).
[9] 2023 Identified small object detection as a major challenge across studies; noted that models perform significantly worse on insects that occupy few pixels, urging focus on multi-scale methods.
[50] 2024 Developed a method for detecting whiteflies on yellow sticky traps (small targets); used high-resolution input and an improved YOLOv5, achieving high precision on pests $\sim$3mm in size.
[51] 2023 Applied image super-resolution on suspected pest regions before classification; demonstrated that upsampling tiny pest images improved identification accuracy in a lab setting.
@@ 8.4 Computational Cost
Reference Year Description
[3] 2025 Emphasized mobile deployment requirements; aimed to develop models that run on smartphones for on-site pest identification.
[7] 2023 Demonstrated a successful on-device pest classification app; used model optimization (quantization, etc.) to achieve real-time inference on mobile hardware.
[23] 2024 Utilized GA-GhostNet (an efficient CNN) to balance accuracy and speed; significantly reduced model parameters while maintaining high precision.
[37] 2024 Designed their hybrid model with edge constraints in mind (e.g., limiting extra parameters from HP-MHA module) to ensure feasibility of deployment.
[52] 2022 Created a lightweight YOLOv3-based detector ($\sim$5M params) that runs $\sim$20 FPS on edge device; achieved >>80% mAP, illustrating a strong accuracy-speed trade-off for field use.
[53] 2024 Improved YOLOv8 architecture for efficiency (small dataset of 7 classes); obtained 82.3% mAP with only 2M parameters, suitable for low-power deployment.
[54] 2015 Proposed knowledge distillation, a method yet to be fully explored in pest classification, where a smaller model learns from a larger model’s outputs, potentially yielding a compact, accurate pest classifier.